a teacher is four times as old as a student. in 12 years, the student's age will be half of the teacher's age. how old are they now
Answer by stanbon(75887)

a teacher is four times as old as a student. in 12 years, the student's age will be half of the teacher's age. how old are they now
-----
Equations:
t = 4s
s+12 = (1/2)(t+4)
----
Substitute for "t" and solve for "s":
s + 12 = (1/2)(4s+4)
s + 12 = 2s + 2
s = 10 (students's age now)
------
Solve for "t":
t = 4s
t = 40 (teacher's age now)
-------
Cheers,
Stan H.
===========